OpenStudy (anonymous):

A bag contains 7 red marbles, 12 orange marbles, 18 yellow marbles, and 5 purple marbles. A marble is drawn at random from the bag. What is the probability that the marble drawn will not be orange? 5/7 , 1/30 , 2/7 , 1/12

OpenStudy (anonymous):

@aum

OpenStudy (aum):

Find total number of marbles. Number of non-orange marbles = Total number of marbles - No. of orange marbles. Probability of picking a non-orange marble = Number of non-orange marbles / Total number of marbles

OpenStudy (anonymous):

total number of marbles is 42. number of non-orange marbles is 31.

OpenStudy (anonymous):

@aum ok, I did that, not im confused

OpenStudy (anonymous):

sorry, I'm terrible at this :/

OpenStudy (aum):

total number of marbles is 42 is correct number of non-orange marbles is 42 - 12 = 30

OpenStudy (anonymous):

oh! okay so would it be 1/30? or no?

OpenStudy (aum):

Probability of picking a non-orange marble = Number of non-orange marbles / Total number of marbles = 30 / 42 = 5 / 7

OpenStudy (anonymous):

thank you so much